OpenStudy (anonymous):

Hi, I need help to solve this problem please. There are 6 rows of 16 chairs set up for the 3 grade play, in the first 4 rows 2 chairs on each end are reserved for teachers, the rest of the chairs are for students. How many chairs are there for students?

undeadknight26 (undeadknight26):

well it would be 6 x 14...

undeadknight26 (undeadknight26):

No srry scratch that...

OpenStudy (emilylauren):

altogether there are 96 chairs, right?

undeadknight26 (undeadknight26):

4 x 14 = 56 2 x 16 = 32 32 + 56 = 88

OpenStudy (emilylauren):

just 16x6=96 the teachers are in the first 4 rows, using 2 chairs on each end. so teachers are using 16 chairs altogether.

undeadknight26 (undeadknight26):

remember for the first 4 rows there is 14 chairs for the students.

OpenStudy (emilylauren):

So then 96-16=80 which is your answer :)

undeadknight26 (undeadknight26):

...

OpenStudy (anonymous):

6x16=96 2x4=8 96-8=88

undeadknight26 (undeadknight26):

what you said :D

undeadknight26 (undeadknight26):

@ranga did i do this right?

OpenStudy (ranga):

6 rows of 16 chairs = 6 * 16 = 96 total number of chairs. In the first 4 rows 2 chairs on each end are reserved for teachers: (2 + 2)*4 = 16 chairs reserved for the teacher (because on the left end 2 chairs and on the right end 2 chairs in the first 4 rows) Chairs left for students = Total chairs - chairs reserved for teachers = 96 - 16 = 80

OpenStudy (anonymous):

thanks for your help and God bless. I'm Sure the answer is 80